The U.S. and Canada both produce aircraft engines and tons of chemicals, which are sold for the same price in both countries. Suppose that with one unit of labor and one unit of capital, the U.S. can produce either 20 engines or 38 tons of chemicals and Canada can produce either 15 engines or 35 tons of chemicals. Which country has a comparative advantage producing engines?
Answer #1

Opportunity cost in the US for producing engines = 38/20 = 1.9

Opportunity cost in Canada for producing engines = 35/15 = 2.33

As the opportunity cost of producing engines is less in the US i.e. they have to give up only 1.9 units of chemical. they will have a comparative advantage in the production of engines.